10ml of 1N HCl is mixed with 20ml of 1M H2SO4 and 30ml of 1M NaOH. The resultant solution has 
  1. 20 meq of H+ ions
  2. 20 meq of OH- ions
  3. 0 meq of H+ or OH-
  4. 30 milli moles of H+
b)
c)
d)
 
  

Grade:11

2 Answers

Sandeep Mishra
4 years ago
n-factor of H2SO4 is 2, so its meq is 20*2*1=40 meq.
meq of HCl is 10 meq
meq of H+ is 50 meq
for NaOH, meq is 30 meq.
net H+ meq is= 50-30=20meq
answer is (A)  20 meq of H+
Vikas TU
4 years ago